What volume of an 18M MgCl2 solution is required to make 197 mL solution of 2.69...

  1. What volume of an 18M MgCl2 solution is required to make 197 mL solution of 2.69 M?
Answer #1

Concentration of MgCl, = 2.69 - L Volume of the solution = 197mL X- L 1000mL =0.197 L Number of moles of MgCl, = 2.69 - -X0.1